OpenStudy (anonymous):

-16a^-3b^-5 ------------- = ? 2a^-4b^2 Simplify the expression using properties of exponents.

OpenStudy (anonymous):

similar to the others you have posted before. Since the numbers are multiplied with each other you can split them apart like this \[\large \frac{ -16a^{-3}}{2a^{-4} }*\frac{ b^{-5} }{ b^2 }\] and solve them individually

OpenStudy (anonymous):

the last part is the easiest so let's start there, \[\frac{ b^{-5} }{ b^2 }=?\]

OpenStudy (anonymous):

-8a^7b^7 ?

OpenStudy (anonymous):

close, but not correct

OpenStudy (anonymous):

Your b looks right.

OpenStudy (anonymous):

-8ab^7 ?

OpenStudy (anonymous):

-3--4 =? -5-2 = ?

OpenStudy (anonymous):

it's either -8a ----- b^7

OpenStudy (anonymous):

or -8 ---- ab^7

OpenStudy (anonymous):

\[\frac{ -8a }{ b^7 } = -8a*b^{-7}\]

OpenStudy (anonymous):

the latter option would be \(-8a^{-1} *b^{-7}\) which is different